OpenStudy (anonymous):

limit definition of the derivative to find f ' (x) for f(x)= x2 + 3x + 1

jimthompson5910 (jim_thompson5910):

first can you tell me what f(x+h) is in this case?

OpenStudy (anonymous):

No idea ...

jimthompson5910 (jim_thompson5910):

f(x) = x^2 + 3x + 1 f(x+h) = (x+h)^2 + 3(x+h) + 1 ... replace all 'x' terms with 'x+h' f(x+h) = x^2 + 2xh + h^2 + 3x + 3h + 1

jimthompson5910 (jim_thompson5910):

now you have this, use this and plug it into [f(x+h) - f(x)]/h

OpenStudy (anonymous):

why do we use f(x+h) ?

OpenStudy (anonymous):

and what am I plugging into [f(x+h) - f(x)]/h ?

jimthompson5910 (jim_thompson5910):

sry the site is being slow

jimthompson5910 (jim_thompson5910):

we use f(x+h) because if you drew this out, you'll have a secant line at (x,f(x)) and (x+h, f(x+h)) a drawing might help see this now imagine dragging one point closer to another fixed point....you'll have that secant line slowly turn into a tangent line...which is exactly what the derivative helps us find

jimthompson5910 (jim_thompson5910):

anyways [f(x+h) - f(x)]/h [( x^2 + 2xh + h^2 + 3x + 3h + 1) - ( x^2 + 3x + 1)]/h [x^2 + 2xh + h^2 + 3x + 3h + 1 - x^2 - 3x - 1]/h [2xh + h^2 + 3h]/h [h(2x + h + 3)]/h 2x + h + 3

jimthompson5910 (jim_thompson5910):

Then you take the limit as h ---> 0 ie you plug in h = 0 to get 2x + h + 3 2x + 0 + 3 2x + 3

jimthompson5910 (jim_thompson5910):

so the derivative function is f ' (x) = 2x + 3

OpenStudy (anonymous):

okay thank you so much ! that makes a lot more sense especially when I graphed it

jimthompson5910 (jim_thompson5910):

yeah a picture says 1000 words often, so I'm glad it's clicking